• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Z.-E. Meziani

    Log entry time 00:10:25 on July24,2003

    Entry number 106386

    keyword=Swing shift summary

    16:00 Swing shift starts
    Waiting for beam...
    18:42 Beam back on Empty target
    Asked MCC to adjust the orbit so that the positions on BPMB and BPMA
    are x = -1+/-0.1 and y= 0.0+/- 0.1
    18:50 Beam off
    19:05 MCC informed us that that have secured A and B and are tuning Hall C
    19:30 Still waiting for beam...
    20:00 Beam back, requested alignment.
    20:10 Alignment Ok on BPMB (A)
    20:12 Moving target to Beo+12C long foil to check position on TV monitor
    Raster off.
    Beam seems to be offset on the Beo by 5mm too high and 5mm to the right.
    We don't know were the target is
    20:40 Performed Harp Scan
    run # 2168
    Harp5 132 um Hor. 150 um Ver.
    Harp6 121 um Hor. 137 um Ver.
    20: 45 Asked for controlled access to check the alignment and
    remove the Beo in Beo+12C long foil
    22:00 Asked for beam back
    22:10 Beam back and requested BPMB and BPMA at x = 0.0 and y =0.0
    Start Optics measurements
    22:18 run#2715 single 12C foil in elastic kin. 1.2M
    22:33 run#2716 multifoil 12C in elastic kin. 1.4M
    22:44 run #2717 empty target in elastic kin. 1.1M
    Performed NMR measurement
    23:07 run #2718 single foil elastic kin. + 3% (bad run)
    Recycled Q2 and Q3
    23:27 run #2719 single foil elastic kin. +3% 1.2 M
    23:42 run#2720 multifoil elastic kin. +3% 1M
    Beam off
    24:00 Shift ends